Graph the line with slope -2 passing through the point (-2,2)
Thanks again for help with graphing

Answer by Alan3354(69443)

Graph the line with slope -2 passing through the point (-2,2)
Thanks again for help with graphing
--------------------
Find the equation.
This is always done the same way:
y-y1 = m*(x-x1) where m is the slope and (x1,y1) is the point
y-2 = -2*(x+2)
y-2 = -2x-4
y = -2x - 2
-----------
The y-intercept is (0,-2). Draw a straight line thru that and the other point.